# Siqi Kou1∗, Jiachun Jin1, Chang Liu1∗, Ye Ma2, Jian Jia2, Quan Chen2, Peng Jiang2, Zhijie Deng1†

### v1

@@ 4.2 An Efficient Strategy for Constructing Orthus
Initialized with the typical Chameleon-7B [53], Orthus can acquire image processing capabilities while preserving the text generation of Chameleon after 9-hour training on 10k high-quality images [19] using 8 A100 GPUs. We designate this model as Orthus-base. Figure 4 shows its image and text (as well as mixed-modality) generation prowess.
Although the decoder in the VQ-VAE [59] used by Chameleon can reconstruct the raw image pixels given the patch-wise features $V$ to some extent, it can be suboptimal due to the quantization-aware training. To address this, we advocate further tuning its decoder to reconstruct high-quality images directly based on $V$. The comparison between the capacity of the original VQ-VAE of Chameleon and ours is exhibited in Appendix A.
@@ 4.3 Multimodal Post-training
Following the typical training procedure of multimodal models [67, 61], we further refine Orthus-base by fine-tuning it on high-quality text-image pairs and instructional data. All parameters except for those of the vision autoencoder are tuned. Following the input format of Chameleon, we surround image features $V$ with the embeddings of special tokens [BOI] and [EOI] before the concatenation with text embeddings. A [SEP] token is used to separate user input and model output in each conversation.
Let $\mathcal{L}_{\text{ar}}$ denote the AR loss on the text tokens. The entire training objective of Orthus is then $\mathcal{L}_{Orthus}=\mathcal{L}_{\text{ar}}+\lambda\mathcal{L}_{\text{diff}}$, where $\lambda$ is a balancing coefficient. Hereinafter, we will denote the model trained following this objective as Orthus, distinguishing it from Orthus-base.
@@ 5 Experiments
In this section, we evaluate Orthus’s performance across multimodal understanding, visual generation, and mix-modality generation tasks. Both quantitative and qualitative results demonstrate the effectiveness of Orthus.
5.1 Experimental Setup
Implementation Details. We implement the diffusion head as an MLP consisting of 3 residual blocks, each sequentially applying AdaLN [38], a linear layer (width of 1536 channels), SiLU activation, and another linear layer. The condition vector ${\bm{f}}_{i}$ is added to the diffusion time embedding, which is then incorporated through AdaLN. The diffusion noise schedule is linear following [44], with 1000 steps at training time. During the training of Orthus-base, we employ AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.99$) with a learning rate of 1e-4 and a cosine decay scheduler. During post-training from Orthus-base to Orthus, we mix LlaVA-v1.5-665K [34] instruction tuning data and high-quality text-to-image data (JourneyDB [48] and LAION-COCO-aesthetic [19] recaptioned from ShareGPT-4v [3]) at a ratio of 1:1 and optimize with a learning rate of 1e-5. $\lambda$ is set to $100$ during post-training. During inference, we use greedy decoding to generate text. For image generation, we adopt the DDIM [46] sampler with 100 steps. We employ classifier-free guidance (CFG) [20] with the scale set to 5 during sampling. All images are generated at a resolution of 512$\times$512. Both training and evaluation are carried out on servers equipped with 8 NVIDIA A100 80GB GPUs.
Baselines. We make comparisons with: (i) image-understanding models including LlaVA [32, 34], InstructBLIP [7], Qwen-VL-Chat [1], mPLUG-Owl2 [71], and IDEFICS-9B [26]; (ii) image-generation models including LlamaGen [49], SDv1.5 [44], PixArt-$\alpha$ [2], SDv2.1 [44], DALL-E 2 [42], SDXL [39], and SD3 [12]; (iii) unified models including Emu [50], NExT-GPT [64], SEED-X [15], Gemini-Nano-1 [54], LWM [35], Transfusion [77], and Show-o [67]. We also fine-tune pre-trained Chameleon [6] with the same high-quality post-training dataset as Orthus to provide an apple-to-apple baseline.
5.2 Multimodal Understanding
Benchmarks. Following LLaVa [34], we evaluate the multimodal understanding capabilities of Orthus on POPE [29], MME [14], VQAv2 [18], GQA [22], and MMMU [75] and report the exact match accuracy.
@@ 5.2 Multimodal Understanding
Und. Only LlaVa [34] 7B 76.3 809.6 - - -
LlaVA-v1.5 [32] 7B 85.9 1510.7 78.5 62.0 35.4
InstructBLIP [7] 7B - - - 49.2 -
Qwen-VL-Chat [1] 7B - 1487.5 78.2 57.5 -
mPLUG-Owl2 [71] 7B 85.8 1450.2 79.4 56.1 -
IDEFICS-9B [26] 8B - - - 50.9 -
InstructBLIP [7] 13B 78.9 1212.8 - 49.5 -
Und. and Gen. $\text{Emu}^{*}$ [50] 13B - - 52.0 - -
@@
Gemini-Nano-1 [54] 1.8B - - 62.7 - 26.3
Show-o [67] 1.3B 73.8 948.4 59.3 48.7 25.1
LWM [35] 7B 75.2 - 55.8 44.8 -
@@
Table 1: Evaluation on multimodal understanding benchmarks. Und. and Gen. denote “understanding” and “generation”, respectively. Models using external pre-trained diffusion models are marked with * and $\text{Chameleon}^{\dagger}$ is post-trained with the same dataset as Orthus. The results in bold and underline are the best and second-best results, respectively. The results correspond to the exact match accuracy.
Results. Table 1 summarizes the capabilities of Orthus on both domain-specific visual question answering and general visual comprehension tasks. (i) Compared to Chameleon post-trained with the same dataset, Orthus consistently demonstrates superior performance across all benchmarks. Besides, inspecting OCR-related tasks in MME-P, we witness a significant superiority of Orthus over Chameleon (with scores of 70 vs. 45). These results validate the superiority of Orthus’s modeling by adopting lossless representations for images. (ii) On the GQA benchmark, Orthus outperforms larger, understanding-only models such as InstructBLIP-13B and IDEFICS-9B. Moreover, it achieves competitive results with models like Qwen-VL-Chat and mPLUG-Owl2, which have been trained extensively on task-specific diverse VQA datasets. We can expect enhanced visual understanding capabilities of Orthus by incorporating more high-quality data for post-training. (iii) Orthus outperforms other unified models using a single transformer like LWM and Show-o across all benchmarks, highlighting its efficacy for unified modeling. (iv) Compared to larger unified models using an external diffusion model and pre-trained with extensive image-text pairs, such as NExT-GPT-13B, Orthus achieves decent results on the VQAv2 benchmark. It is reasonable to speculate that Orthus’s potential for multimodal understanding problems can be further unleashed by scaling up training compute and data.
@@
5.3 Visual Generation
Benchmarks. Following SD3 [12], we evaluate Orthus’s visual generation performance on GenEval [16], a benchmark that examines the compositional capabilities of the image generation model through a detailed instance-level analysis. Additionally, we evaluate on HPSv2 [65], a benchmark based on human preferences.
Quantitative Results. We report GenEval scores across six dimensions and the HPSv2 score in Table 2. (i) When compared with strong competitors specialized for text-to-image generations such as DALL·E 2 and SDXL, Orthus achieves an improvement of 0.06 and 0.03 on GenEval, respectively. While Orthus lags behind SD3, it’s worth noting that SD3 benefits from synthetic image captions generated through backtranslation, whereas Orthus does not employ this trick. (ii) Compared to Chameleon and its post-trained version, Orthus demonstrates significant superiority on both GenEval and HPSv2. This advantage can be attributed to the utilization of continuous image representations and diffusion-based continuous modeling, which facilitates the generation of high-quality images with richer detail and stronger alignment with human preferences. (iii) Compared with other unified models such as SEED-X, LWM, and Show-o, Orthus obtains significantly better performance, highlighting the advantages of its modeling strategy. Orthus also achieves comparable results to Transfusion, which leverages a larger dataset of text-image pairs during pretraining. (iv) Orthus outperforms both SDv1.5 and SDv2 in HPSv2, showing its ability to generate high-quality, aesthetically pleasing images aligned with human preference.
Qualitative results. Figure 3 showcases images generated by Orthus alongside results from other unified models, including Chameleon and Show-o. Results show that Orthus is capable of generating diverse, engaging, and realistic visual imagery at the resolution of 512$\times$512.
Type Model Res. GenEval $\uparrow$ HPSv2$\uparrow$
Single Obj. Two Obj. Count. Colors Position Color Attri. Overall
Gen. Only LlamaGen [49] 512 0.71 0.34 0.21 0.58 0.07 0.04 0.32 26.3
SDv1.5 [44] 512 0.97 0.38 0.35 0.76 0.04 0.06 0.43 27.0
PixArt-$\alpha$ [2] 512 0.98 0.50 0.44 0.80 0.08 0.07 0.48 -
SDv2.1 [44] 512 0.98 0.51 0.44 0.85 0.07 0.17 0.50 27.2
DALL-E 2 [42] 512 0.94 0.66 0.49 0.77 0.10 0.19 0.52 26.9
SDXL [39] 512 0.98 0.74 0.39 0.85 0.15 0.23 0.55 30.9
SD3(d=30) [12] 512 0.96 0.80 0.65 0.73 0.33 0.37 0.64 -
[…]
Chameleon [53] 512 - - - - - - 0.39 -
LWM [35] 256 0.93 0.41 0.46 0.79 0.09 0.15 0.47 26.1
Show-o [67] 256 0.95 0.52 0.49 0.82 0.11 0.28 0.53 27.3
Transfusion [77] 256 - - - - - - 0.63 -
$\text{Chameleon}^{\dagger}$ 512 0.98 0.47 0.18 0.75 0.06 0.15 0.43 26.9
Orthus (Ours) 512 0.99 0.75 0.26 0.84 0.28 0.38 0.58 28.2
Table 2: Comparison with state-of-the-arts on visual generation benchmarks. Und. and Gen. denote “understanding” and “generation”, respectively. Model using external pre-trained diffusion model is marked with * and $\text{Chameleon}^{\dagger}$ is post-trained with the same dataset as Orthus. The results in bold and underline are the best and second-best results, respectively.
5.4 Mixed-modality Generation
In this section, we investigate the mixed-modality generation capabilities of Orthus. We first show that Orthus-base can perform mixed-modality generation, as demonstrated in the left of Figure 4. Orthus-base generates a coherent sequence of interleaved image-text segments based on instructions from users in a zero-shot manner.
To further validate Orthus’s superiority in modeling interleaved data, we fine-tune Orthus-base with the unified learning objective on the StoryStream [69] dataset, which includes a collection of images and corresponding narratives from cartoon series. As shown in Figure 4, after training, Orthus can generate contextually consistent scenes paired with narrative text given an initial image-text pair and the instruction “Please continue this story.” Notably, there is a strong alignment between images and text (e.g., the smile on the monkey’s face) as well as consistent detail across images (e.g., the boy’s orange hat). These results highlight Orthus’s potential for various applications, including report generation, educational content creation, and others requiring seamless multimodal content generation.
Figure 4: Qualitative results of interleaved image-text generation. Left: Zero-shot results generated by Orthus-base which is only trained for 72 A100 GPU hours. Right: Interleaved storybook creation results after finetuned on the StoryStream [69] dataset. Results show that Orthus excels in generating logically coherent interleaved image-text with high relevance.
@@ 6 Conclusion
Supplementary Material
@@ Appendix A Comparison of Vision Autoencoder
To construct a vision autoencoder capable of decoding high-quality images based on continuous image features $V$, we freeze the encoder of Chameleon’s VQ-VAE, drop the quantization step, and finetune the decoder only to reconstruct images, transforming it into a conventional continuous autoencoder effectively. Table 5 shows that our vision autoencoder achieves better reconstruction quality compared to the original VQ-VAE. The evaluation is on the subset of LAION-Aesthetic [19] containing 10000 images.
@@
Appendix B Detailed Scores of HPSv2
Table  6 reports scores of HPSv2 [65] across four categories: ‘Anime’, ‘Concept art’, ‘Paintings’, and ‘Photo’.
[…]
LDM [44] 25.7 25.2 25.3 27.0 25.8
DALL-E 2 [42] 27.3 26.5 26.7 27.2 26.9
SDv1.5 [44] 27.3 26.6 26.7 27.3 27.0
SDv2 [44] 27.5 26.9 26.9 27.5 27.2
SDXL [39] 36.0 34.9 35.7 31.9 30.9
Show-o [67] 27.9 27.3 27.1 27.0 27.3
$\text{Chameleon}^{\dagger}$ 26.9 27.2 27.3 26.0 26.9
Orthus(Ours) 28.2 28.4 28.4 27.8 28.2
Table 6: Comparison of state-of-art baselines on HPSv2.
Appendix C Importance of Diffusion Modeling
@@
Appendix D Qualitative Results on Visual Understanding
In addition to quantitatively evaluating Orthus on domain-specific tasks in Section 5.2, we also assess its performance in general chat scenarios. Results in Figure 6 demonstrate that Orthus exhibits strong instruction-following capabilities and robust generalization abilities.
Figure 6: Qualitative analysis of Orthus on visual understanding.
Appendix E Qualitative Results on Visual Generation
Figure 7 shows examples of images generated from Orthus.

### v2

@@ 4.2 An Efficient Strategy for Constructing Orthus-base
Initialized with the typical Chameleon-7B Team (2024), Orthus can acquire image processing capabilities while preserving the text generation capacity after 9-hour training on 10k high-quality images laion-coco aesthetic using 8 A100 GPUs. We designate this model as Orthus-base, a pre-trained model capable of generating continuous image features and discrete text tokens.
Although the decoder in the VQ-VAE Van Den Oord et al. (2017) used by Chameleon can reconstruct the raw image pixels given the patch-wise features $V$ to some extent, it can be suboptimal due to the quantization-aware training. To address this, we advocate further tuning its decoder to reconstruct high-quality images directly based on $V$. The comparison between the capacity of the original VQ-VAE and ours is exhibited in Appendix A.
@@ 4.3 Multimodal Post-training
Orthus-base can be further post-trained to unlock its potential for interleaved image-text modeling in complex downstream tasks. These include generating text (e.g., visual question answering), images (e.g., image editing), or even both images and text (e.g., storybook generation) from mixed-modality inputs. For example, given $[V,U]$ as user input and $[V,U,V,U]$ as model output, we surround image features $V$ with the embeddings of special tokens [BOI] and [EOI] before the concatenation with $U$. A [SEP] token is used to separate user input and model output in each conversation.
Let $\mathcal{L}_{\text{ar}}$ denote the AR loss on the text tokens. The entire training objective of Orthus is then $\mathcal{L}_{Orthus}=\mathcal{L}_{\text{ar}}+\lambda\mathcal{L}_{\text{diff}}$, where $\lambda$ is a balancing coefficient. All parameters except for those of the vision autoencoder are tuned. Hereinafter, we will denote the model trained following this objective as Orthus, distinguishing it from Orthus-base.
@@ 5 Experiments
In this section, we evaluate Orthus’s performance in interleaved image-text modeling as well as visual understanding and generation. Both quantitative and qualitative results demonstrate the effectiveness of Orthus.
5.1 Implementation Details
We implement the diffusion head as an MLP consisting of 3 residual blocks, each sequentially applying AdaLN Peebles & Xie (2023), a linear layer (width of 1536 channels), SiLU activation, and another linear layer. The condition vector ${\bm{f}}_{i}$ is added to the diffusion time embedding, which is then incorporated through AdaLN. The diffusion noise schedule is linear following Rombach et al. (2022), with 1000 steps at training time. $\lambda$ is set to $100$ to balance the order of magnitude between $\mathcal{L_{\text{diff}}}$ and $\mathcal{L_{\text{ar}}}$ during post-training. During inference, we use greedy decoding to generate text. For image generation, we adopt the DDIM Song et al. (2020a) sampler with 100 steps. We employ classifier-free guidance (CFG) Ho & Salimans (2022) with the scale set to 5 during sampling. All images are generated at a resolution of 512$\times$512. More training and evaluation details are provided in Appendix B.
5.2 Interleaved Image-Text Generation
Figure 3: Qualitative results on mixed image-text understanding and generation of Orthus. Left: Image editing results after fine-tuned on Instruct-Pix2Pix Brooks et al. (2023). Notably, Orthus exhibits in-context learning capacity by performing image editing successfully when provided with examples rather than explicit instructions, which is not included in the training dataset. Right: Interleaved storybook creation results after finetuned on the StoryStream Yang et al. (2024b) dataset. Results show that Orthus excels in generating logically coherent interleaved image-text with high relevance.
Compared to existing unified models, such as Janus-series Wu et al. (2024a); Ma et al. (2024), that focus exclusively on visual understanding and generation, we investigate Orthus’s flexibility and extensibility to model interleaved images and text on two representative downstream tasks: image editing and storybook generation.
Table 1: Comparisons of CLIP similarities Ruiz et al. (2023); Gal et al. (2022) between editing-specific diffusion models and Orthus on the test dataset of Instruct-Pix2Pix.
@@ 5.2 Interleaved Image-Text Generation
PnP Tumanyan et al. (2023) 0.156 0.76 0.023
SDEdit Meng et al. 0.229 0.84 0.047
I-Pix2Pix Brooks et al. (2023) 0.233 0.88 0.045
Orthus (Ours) 0.238 0.87 0.049
Image-Text $\bm{\rightarrow}$ Image. We compare the performance of Orthus with an editing-specific diffusion model after training Orthus-base on the 400k Instruct-pix2pix Brooks et al. (2023) training dataset. Table 1 shows that the images edited by Orthus align well with both the given instruction and the input image, performing comparably to or even surpassing the editing-specific diffusion model. Moreover, Figure 3 illustrates Orthus’s strong generalization ability to edit images in zero-shot real-image domains. Notably, Orthus exhibits in-context learning capacity as a unified multimodal model: when provided with examples instead of explicit instructions that do not match the formats seen during training, it successfully completes the task. This highlights Orthus’s strong capability for interleaved data modeling and its great potential as a foundation multimodal model. More editing examples and comparisons are provided in Appendix F.
Image-Text $\bm{\rightarrow}$ Image-Text-Image-Text. To further validate Orthus’s superiority in modeling interleaved data, we fine-tune Orthus-base with the unified learning objective on the StoryStream Yang et al. (2024b) dataset, which includes a collection of images and corresponding narratives from cartoon series. As shown in Figure 3, after training, Orthus can generate contextually consistent scenes paired with narrative text given an initial image-text pair and the instruction “Please continue this story.” Notably, there is a strong alignment between images and text (e.g., the smile on the monkey’s face) as well as consistent detail across images (e.g., the boy’s orange hat and red scarf). These results highlight Orthus’s ability to generate long sequences of contextually relevant images and text. This capability opens up potential applications such as report generation, educational content creation, and other tasks requiring seamless mixed-modality content generation.
Table 2: Evaluation on visual understanding benchmarks. Und. and Gen. denote “understanding” and “generation”, respectively. Models using external pre-trained diffusion models are marked with * and $\text{Chameleon}^{\dagger}$ is post-trained with the same dataset as Orthus. The results in bold and underline are the best and second-best results, respectively. The results correspond to the exact match accuracy.
@@
Und. Only LlaVa Liu et al. (2024d) 7B 76.3 809.6 - - -
LlaVA-v1.5 Liu et al. (2024b) 7B 85.9 1510.7 78.5 62.0 35.4
InstructBLIP Dai et al. (2023) 7B - - - 49.2 -
Qwen-VL-Chat Bai et al. (2023) 7B - 1487.5 78.2 57.5 -
Emu3-Chat Wang et al. (2024) 8B 85.2 1243.8 75.1 60.3 31.6
InstructBLIP Dai et al. (2023) 13B 78.9 1212.8 - 49.5 -
Und. and Gen. $\text{Emu}^{*}$ Sun et al. (2023) 13B - - 52.0 - -
[…]
Gemini-Nano-1 Team et al. (2023) 1.8B - - 62.7 - 26.3
Show-o Xie et al. (2024) 1.3B 73.8 948.4 59.3 48.7 25.1
LWM Liu et al. (2024e) 7B 75.2 - 55.8 44.8 -
@@
5.3 Visual Understanding and Generation
In this section, we validate the effectiveness of Orthus on visual understanding and generation by post-training Orthus-base with a mixture of LlaVA-v1.5-665K Liu et al. (2024d) and high-quality text-to-image data (JourneyDB Sun et al. (2024a) and LAION-COCO-aesthetic laion-coco aesthetic recaptioned from ShareGPT-4v Chen et al. (2023a)). We also fine-tune pre-trained Chameleon Chern et al. (2024) with the same mixed dataset as Orthus to provide an apple-to-apple baseline.
Image $\bm{\rightarrow}$ Text. Table 2 shows that: (i) Compared to Chameleon post-trained with the same dataset, Orthus consistently demonstrates superior performance across all benchmarks. Besides, inspecting OCR-related tasks in MME-P, we witness a significant superiority of Orthus over Chameleon (with scores of 70 vs. 45). These results validate the superiority of Orthus’s modeling by adopting lossless representations for images. (ii) Orthus outperforms other unified models using a single transformer like LWM and Show-o across all benchmarks, highlighting its efficacy for unified modeling. (iii) Compared to larger unified models using an external diffusion model, such as NExT-GPT-13B, Orthus achieves decent results on the VQAv2 benchmark. It is reasonable to speculate that Orthus’s potential for multimodal understanding problems can be further unleashed by scaling up training compute and data.
Text $\bm{\rightarrow}$ image.
Table 3: Comparison with state-of-the-arts on visual generation benchmarks. Model using external pre-trained diffusion model is marked with * and $\text{Chameleon}^{\dagger}$ is post-trained with the same dataset as Orthus. The results in bold and underline are the best and second-best results, respectively.
Type Model Res. GenEval HPS
Gen. Only SDv1.5 Rombach et al. (2022) 512 0.43 27.0
SDv2.1 Rombach et al. (2022) 512 0.50 27.2
DALL-E Ramesh et al. (2022) 512 0.52 26.9
Emu3-Gen Wang et al. (2024) 512 0.54 -
SDXL Podell et al. (2023) 512 0.55 30.9
SD3(d=30) Esser et al. (2024) 512 0.64 -
[…]
LWM Liu et al. (2024e) 256 0.47 26.1
Show-o Xie et al. (2024) 256 0.53 27.3
Transfusion Zhou et al. (2024) 256 0.63 -
$\text{Chameleon}^{\dagger}$ 512 0.43 26.9
Orthus (Ours) 512 0.58 28.2
Table 3 shows that: (i) When compared with strong competitors specialized for text-to-image generations such as DALL-E 2 and SDXL, Orthus achieves an improvement of 0.06 and 0.03 on GenEval, respectively. (ii) Compared to Chameleon and its post-trained version, Orthus demonstrates significant superiority on both GenEval and HPSv2. This advantage can be attributed to the utilization of continuous image representations and diffusion-based continuous modeling, which facilitates the generation of high-quality images with richer detail and stronger alignment with human preferences. (iii) Compared with other unified models such as SEED-X, LWM, and Show-o, Orthus obtains significantly better performance, highlighting the advantages of its modeling strategy. (iv) Qualitative results in Figure 4 showcases images generated by Orthus alongside results from other unified models, including Chameleon and Show-o. Results show that Orthus is capable of generating diverse, engaging, and realistic visual imagery at the resolution of 512$\times$512.
@@ 6 Conclusion
Impact Statement
This work presents a challenge in machine learning and proposes a solution, the potential negative consequences are not apparent. While it is theoretically possible for any technique to be misused, the likelihood of such misuse occurring at the current stage is low.
@@ Appendix A Comparison of Vision Autoencoder
To construct a vision autoencoder capable of decoding high-quality images based on continuous image features $V$, we freeze the encoder of Chameleon’s VQ-VAE, drop the quantization step, and finetune the decoder only to reconstruct images, transforming it into a conventional continuous autoencoder effectively. The decoder is trained on LAION-Aesthetic dataset laion-coco aesthetic using a learning rate of 1e-5, a batch size of 256, and a total of 15,000 training steps. Table 6 shows that our vision autoencoder achieves better reconstruction quality compared to the original VQ-VAE. The evaluation is conducted on a subset of the LAION-Aesthetic, consisting of 10,000 images that are excluded from the training dataset.
@@
Appendix B Training details
The images for training Orthus-base are the first 10k from laion-coco aesthetic . Both training and evaluation are carried out on servers equipped with 8 NVIDIA A100 80GB GPUs.
Table 7: Training details for constructing Orthus-base and the instruction-tuned one for visual understanding and generation in 5.3.
      Model       Orthus-base       Instruct-tuning
      Optimizer       AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.99$)
      Learning Rate       1e-4       1e-5
      Batch Size       32       16
      Training Steps       15,000       35,000
Appendix C Diffusion Loss v.s. MSE Loss
@@ Appendix B Training details
Appendix D Examples on Visual Generation
Figure 6 shows examples of images generated from Orthus post-trained in Section 5.3.
@@ Appendix D Examples on Visual Generation
Appendix E Examples on Visual Understanding
In addition to quantitatively evaluating Orthus in Section 5.3 on domain-specific tasks, we also assess its performance in general chat scenarios in Figure 7.
Figure 7: Examples of Orthus on visual understanding. Results demonstrate that Orthus exhibits strong instruction-following capabilities and robust generalization abilities.
Appendix F Examples on Image Editing
Figure 8 shows random examples of image editing by Orthus-base post-trained on Instruct-Pix2Pix Brooks et al. (2023).
Figure 8: Examples of Orthus on image editing. Compared to editing-specific diffusion models Brooks et al. (2023), Orthus demonstrates better fidelity to the original image in regions where no editing is required.